# How Do Large Language Models Understand Graph Patterns? A Benchmark for Graph Pattern Comprehension

### v1

@@ Abstract
Benchmarking the capabilities and limitations of large language models (LLMs) in graph-related tasks is becoming an increasingly popular and crucial area of research. Recent studies have shown that LLMs exhibit a preliminary ability to understand graph structures and node features. However, the potential of LLMs in graph pattern mining remains largely unexplored. This is a key component in fields such as computational chemistry, biology, and social network analysis. To bridge this gap, this work introduces a comprehensive benchmark to assess LLMs’ capabilities in graph pattern tasks. We have developed a benchmark that evaluates whether LLMs can understand graph patterns based on either terminological or topological descriptions. Additionally, our benchmark tests the LLMs’ capacity to autonomously discover graph patterns from data. The benchmark encompasses both synthetic and real datasets, and a variety of models, with a total of 11 tasks and 7 models. Our experimental framework is designed for easy expansion to accommodate new models and datasets. Our findings reveal that: (1) LLMs have preliminary abilities to understand graph patterns, with O1-mini outperforming in the majority of tasks; (2) Formatting input data to align with the knowledge acquired during pretraining can enhance performance; (3) The strategies employed by LLMs may differ from those used in conventional algorithms.
@@ 1 Introduction
Originally trained on textual data, LLMs have demonstrated remarkable success in various tasks, such as reading comprehension and text reasoning (Achiam et al., 2023; Touvron et al., 2023). To evaluate whether LLMs can adapt the text understanding ability across graphs, several studies have investigated this at both the feature and structural levels (Zhao et al., 2023; Chai et al., 2023). Specifically, LLMs have been shown to enhance node features in social networks (Ye et al., 2023; Huang et al., 2024). Additionally, graph structure understanding tasks, such as shortest path and connectivity, have also been evaluated (Guo et al., 2023; Wang et al., 2024).
@@ 3.1 Pattern Translation
Mixtral 1.00 0.00 0.00 0.00 0.00 0.00 1.00 0.25 0.00 0.00 0.00 0.00 1.00 0.11 0.02 0.08 1.00 0.04 0.45 0.05
Gemini 0.52 0.82 0.44 0.32 0.24 0.90 0.84 0.88 0.50 0.23 0.78 0.86 0.90 0.89 0.90 0.78 0.94 0.92 0.67 0.73
Claude 1.00 0.61 0.02 0.00 0.24 0.67 0.78 0.58 0.88 0.62 0.20 0.93 1.00 0.29 0.90 0.53 0.84 0.59 0.65 0.54
Llama 1.00 0.48 0.00 0.00 0.86 0.75 1.00 0.80 0.14 0.30 0.00 0.00 1.00 0.90 0.98 0.77 1.00 0.00 0.66 0.44
GPT-4 1.00 0.12 0.35 0.75 0.98 0.33 0.92 0.76 0.86 0.80 0.42 0.94 1.00 0.28 0.98 0.45 1.00 0.54 0.83 0.55
GPT-4o 0.98 0.29 0.92 0.46 0.74 0.80 0.82 0.77 0.68 0.40 0.88 0.26 1.00 0.64 0.82 0.87 0.68 0.77 0.84 0.58
O1-mini 1.00 0.79 0.62 0.70 1.00 0.94 0.74 0.87 0.66 0.97 0.82 0.90 0.98 0.75 0.96 0.82 0.98 0.85 0.86 0.84
AVG. 0.93 0.44 0.34 0.32 0.58 0.63 0.87 0.70 0.53 0.47 0.44 0.56 0.98 0.55 0.79 0.61 0.92 0.53 0.71 0.54
@@ 4 Topology-based Pattern
4.1 Pattern Isomorphic Mapping
The results, presented in Table 4, show that most LLMs perform well on small datasets, but their performance degrades as the graph scale increases. However, GPT-4o and Claude still maintain high accuracy using the edge list format. Consequently, we select medium- and large-scale graphs for further analysis. Upon manually reviewing the output text from GPT-4o, Claude, and O1-mini, we found that there exist two algorithms to help LLMs reason whether the graphs are isomorphic. One approach involves first calculating node degrees and then mapping nodes with matching degrees, while the other compares the edge list directly. Due to the LLMs’ inherent difficulty with counting, results based on degree counting tend to be less reliable. As shown in Figure 2, O1-mini uses the degree counting approach with the proportion 89%. Only about 30% of the outputs from Claude and GPT-4o use degree counting, with the majority relying on edge comparison. As LLMs are not good at counting, the overall accuracy of Claude and GPT-4o is significantly higher than that of O1-mini.
@@ 5.1 Dense subgraph mining (k-core)
5.2 Frequent Subgraph Extraction
Mining frequent subgraphs is an important task on graphs, defined as finding subgraphs that appear frequently in a graph dataset given a frequency threshold. In this task, we set a 100 turns evaluation. In each turn, we randomly select 10 graphs, and task LLMs to extract frequent patterns and output patterns in topology-based description. The accuracy is calculated when the patterns appear in more than 60% graphs, and summarize the results in Table 7.
@@ 5.2 Frequent Subgraph Extraction
LLMs exhibit a strong capability in identifying frequent subgraphs, with GPT-4 and O1-mini showing impressive performance. However, LLMs are prone to detect simpler patterns rather than more complex ones in the dataset. In Figure 5, we illustrate the accuracy when different models extract the defined patterns in the datasets. Although various patterns appear with similar probability across different datasets, LLMs predominantly identify triangles. The house pattern, a combination of a triangle and square, is rarely recognized.
@@ 5.3 Discriminative Pattern Learning
The results in Table 8 demonstrate that most LLMs successfully perform graph classification tasks, particularly when using the edge list format. Interestingly, a high D.P. score does not always correspond to high accuracy. For instance, despite having a D.P. score of 0.08, O1-mini still achieves a classification accuracy of 0.72.
@@ 6 Evaluation on Real-world Graphs
In this section, we gather a wide range of real-world datasets across diverse domains, including molecules, social networks, bioinformatics, and computer vision. Different from synthetic dataset, the real-world graphs have node attributes and potential noise in the graph structures, which makes the graph pattern tasks more challenging. More details of datasets and prompts are provided in the appendix C. In this evaluation, we focus on pattern detection and discriminative pattern learning.
@@ 7 Discussions
In this section, we endeavor to provide intuitions and insights on LLMs’ ability in graph pattern tasks. We first provide our theoretical intuitions from the perspective of expressiveness and then summarize key insights from our observations.
@@ 7.1 Theoretical Intuitions
In this subsection, we discuss the theoretical intuitions of LLMs’ capabilities in graph pattern recognition from an expressiveness standpoint, often used in analyzing Transformers (Feng et al., 2024; Li et al., 2024; de Luca & Fountoulakis, 2024). We demonstrate that Transformers, given graph inputs, can be configured to perform graph pattern tasks. The LOCAL framework Angluin (1980) is a distributed computing paradigm, implemented by the message passing among the nodes. Many graph pattern algorithms such as pattern detection can be efficiently implemented by LOCAL  (Drucker et al., 2014; Korhonen & Rybicki, 2017). Intuitively if Transformers can simulate any LOCAL algorithms, it has the capability for graph pattern tasks. We build the following theorem for this intuition.
@@
The detailed proof is given in Appendix B. The proof intuition is that the attention mechanism is a message-passing between the tokens and each token is executed in a parallel manner. Consequently, this mechanism enables Transformers to simulate the message passing between nodes. The above theorem suggests the existence of such weights.
@@ 7.2 Empirical Insights
LLMs may employ strategies differently from traditional algorithms. Previous theoretical works (Feng et al., 2024; Wu et al., 2024; Li et al., 2024; de Luca & Fountoulakis, 2024) demonstrate that LLMs can perform graph-related tasks by simulating traditional graph algorithms (with chain-of-thought). However, in many cases, the LLMs pretrained by language could use strategies differently from traditional algorithms. First, while frequent subgraph extraction is computationally more challenging than k-core detection, LLMs perform better on the former. This is because LLMs struggle with simple degree counting and fail to execute specific algorithms effectively. Second, the format of the adjacency list and edge list can be easily converted to each other in linear time, but the adjacency list is often better than the edge list in experiments. We hypothesize that the adjacency list format provides a more concise representation of a node’s edges, which aligns well with the locality of attention mechanisms in LLMs (Xiao et al., 2023). Third, in the pattern isomorphic mapping task, O1-mini prefers the degree counting approach with the proportion 89% while only about 30% of the outputs from Claude and GPT-4o use degree counting, as shown in Figure 2.
Input format that aligns with the pretrained knowledge improves the performance. Our experiments demonstrate that LLMs have basic knowledge about graph patterns and they help LLMs in graph pattern tasks. First, in Section 3.1, LLMs can translate the terminologies to the topology descriptions, which shows that LLMs have the basic knowledge about graph patterns. Second, comparing the results in Section 3.3 and those in Section 4.3, we find that terminology-based graph pattern detection is usually better than topology-based ones. It indicates that with terminology as input, LLMs make use of their internal knowledge to improve performance. Third, comparing the results in Section 3 and Section 4, we find that the terminology-based description is often better than topology-based. In most cases, the adjacency list input is better than the edge list input.
@@ 8 Related work
Large Language Models for Graphs. Here we introduce previous work using LLMs based on the types of graph tasks they focused on. The first category is to solve graph algorithm tasks, such as connectivity, degree counting, cycle check, shortest path, and Hamilton path. NLGraph (Wang et al., 2024) and Talk like a graph (Fatemi et al., 2023) translated graph structures into natural language, and leveraged techniques like few-shot prompting or chain-of-thought (CoT) to feed these descriptions directly into LLMs for inference. Since exact computational algorithms can be employed to obtain an accurate reasoning process and results, GraphInstruct (Luo et al., 2024) built the instruction-tuning dataset and fine-tuned LLMs by Lora. Moreover, Graphwiz (Chen et al., 2024a) extended the scale and diversity of the dataset and applied Direct Preference Optimization (DPO) loss to enhance the performance. Recently, GraphToken (Perozzi et al., 2024) utilized a GNN-like encoder to process the input graph and train this encoder with a frozen LLM.
@@
Graph Pattern Discovery. Graph patterns (also known as substructures, graphlets, motifs and subgraphs) have been extensively studied before the era of LLMs due to their crucial role in real-world applications, particularly in chemistry (Murray & Rees, 2009), biology (Hu et al., 2005), and social science (Wu et al., 2022). Numerous exact algorithms have been proposed for pattern discovery. For example, the algorithms designed to identify and count dense pre-defined patterns, such as k-core, k-cycle, and k-clique, were introduced in  Milo et al. (2002); Zhang & Parthasarathy (2012). These methods were further extended to attributed graphs to capture more realistic and complex patterns (Hu et al., 2019). Frequent subgraph mining algorithms were developed to efficiently search for possible patterns (Yan et al., 2008; Elseidy et al., 2014), aiming to reduce the search space as much as possible.
In recent years, deep learning, particularly graph neural networks (GNNs), has become prominent in this area. Many studies (Chen et al., 2020; Zhang et al., 2023; Luo et al., 2022) investigated the theoretical guarantees of specific GNNs capable of identifying certain substructures, and developed subgraph-aware GNNs with strong expressiveness to improve the accuracy of graph learning tasks.
@@ 9 Conclusion
Appendix A Appendix
Appendix B Proof for Theorem 1
@@
Appendix C Dataset
C.1 Primitive graph patterns
We select primitive graph patterns with varying node counts and edge numbers. First, 3-node patterns are the simplest structures in graphs, so we begin with patterns like the triangle, V-structure (V-S), feedforward loop (FFL), and feedback loop (FBL). In directed graphs, the V-S has two edges, while both the FFL and FBL have three. However, the FFL and FBL differ in the direction of one edge. For 4-node patterns, we select the tailed-triangle (T-triangle), square, and diamond for undirected graphs, and the directed-diamond for directed graphs. Both the T-triangle and square have four edges but differ in their connectivity, while the diamond includes five edges. Finally, we introduce the house pattern, a 5-node structure combining a triangle and a square. The summarization of patterns shown in Table 11
@@ C.1 Primitive graph patterns
C.2 Synthetic dataset details
@@
C.3 Real-world dataset details

### v2

@@ Abstract
Benchmarking the capabilities and limitations of large language models (LLMs) in graph-related tasks is becoming an increasingly popular and crucial area of research. Recent studies have shown that LLMs exhibit a preliminary ability to understand graph structures and node features. However, the potential of LLMs in graph pattern mining remains largely unexplored. This is a key component in fields such as computational chemistry, biology, and social network analysis. To bridge this gap, this work introduces a comprehensive benchmark to assess LLMs’ capabilities in graph pattern tasks. We have developed a benchmark that evaluates whether LLMs can understand graph patterns based on either terminological or topological descriptions. Additionally, our benchmark tests the LLMs’ capacity to autonomously discover graph patterns from data. The benchmark encompasses both synthetic and real datasets, and a variety of models, with a total of 11 tasks and 7 models. Our experimental framework is designed for easy expansion to accommodate new models and datasets. Our findings reveal that: (1) LLMs have preliminary abilities to understand graph patterns, with O1-mini outperforming in the majority of tasks; (2) Formatting input graph data to align with the knowledge acquired during pretraining can enhance performance; (3) LLMs employ diverse potential algorithms to solve one task, with performance varying based on their execution capabilities. Our dataset and implementations are available at https://github.com/DDigimon/GraphPattern.
00footnotetext: Correspondence to: Caihua Shan (caihuashan@microsoft.com; Jiliang Tang (tangjili@msu.edu).
@@ 1 Introduction
Originally trained on textual data, LLMs have demonstrated remarkable success in various tasks, such as reading comprehension and text reasoning (Achiam et al., 2023; Touvron et al., 2023). To evaluate whether LLMs can adapt the text understanding ability across graphs, several studies have investigated this at both the feature and structural levels (Zhao et al., 2023; Chai et al., 2023). Specifically, LLMs have been shown to enhance node features on graphs (Ye et al., 2023; Huang et al., 2024; Chen et al., 2024c). Additionally, graph structure understanding tasks, such as shortest path and connectivity, have also been evaluated (Guo et al., 2023; Wang et al., 2024).
@@ 3.1 Pattern Translation
Mixtral 1.00 0.00 0.00 0.00 0.00 0.00 1.00 0.25 0.00 0.00 0.00 0.00 1.00 0.11 0.02 0.00 1.00 0.04 0.45 0.04
Gemini 0.52 0.82 0.46 0.32 0.24 0.90 0.84 0.88 0.50 0.23 0.78 0.76 0.90 0.89 0.90 0.73 0.94 0.91 0.67 0.71
Claude 1.00 0.61 0.02 0.00 0.24 0.67 0.78 0.58 0.88 0.62 0.20 0.53 1.00 0.29 0.90 0.42 0.84 0.42 0.65 0.46
Llama 1.00 0.48 0.00 0.00 0.86 0.75 1.00 0.80 0.12 0.80 0.00 0.00 1.00 0.90 0.98 0.77 1.00 0.00 0.66 0.50
GPT-4 1.00 0.12 0.35 0.75 0.98 0.33 0.92 0.76 0.86 0.80 0.42 0.80 1.00 0.28 0.98 0.43 1.00 0.54 0.83 0.53
GPT-4o 0.98 0.29 0.92 0.46 0.74 0.80 0.82 0.77 0.68 0.41 0.88 0.05 1.00 0.64 0.82 0.81 0.68 0.54 0.84 0.53
O1-mini 1.00 0.79 0.62 0.35 1.00 0.80 0.74 0.87 0.66 0.95 0.82 0.85 0.98 0.75 0.96 0.80 0.98 0.84 0.86 0.78
AVG. 0.93 0.44 0.34 0.32 0.27 0.63 0.87 0.69 0.53 0.54 0.44 0.49 0.98 0.55 0.79 0.57 0.92 0.47 0.71 0.51
@@ 4 Topology-based Pattern
4.1 Graph Isomorphic Mapping
The results, presented in Table 4, show that most LLMs perform well on small datasets, but their performance degrades as the graph scale increases. However, GPT-4o and Claude still maintain high accuracy using the edge list format. Consequently, we select medium- and large-scale graphs for further analysis. Upon manually reviewing the output text from GPT-4o, Claude, and O1-mini, we found that there exist two algorithms to help LLMs reason whether the graphs are isomorphic. One approach involves first calculating node degrees and then mapping nodes with matching degrees, while the other compares the edge list directly. Due to the LLMs’ inherent difficulty with counting, results based on degree counting tend to be less reliable. As shown in Figure 2, O1-mini uses the degree counting approach with the proportion 89%. Only about 30% of the outputs from Claude and GPT-4o use degree counting, with the majority relying on edge comparison. As LLMs are not good at counting, the accuracy of Claude and GPT-4o is significantly higher than O1-mini.
@@ 5.1 Dense subgraph mining (k-core)
5.2 Frequent Subgraph Extraction
Mining frequent subgraphs is an important task on graphs, defined as finding subgraphs that appear frequently in a graph dataset given a frequency threshold. For each pattern, we first generate a graph dataset, ensuring that each graph contains the target pattern. The statistics of datasets are provided in Table 12. In each turn, we randomly select 10 graphs from the dataset, task LLMs to extract frequent patterns based on these selected graphs, and output patterns in the topology-based description. We repeat this process 100 times and calculate the accuracy as the percentage of cases where the output pattern appears in more than 60% of the selected graphs. It is worth noting that the extracted pattern does not need to exactly match the tested pattern. For example, if the LLMs identify a triangle pattern during testing with a house pattern, we still consider this an accurate outcome. The accuracy and frequency of extracted patterns are summarized in Table 7 and Figure 5, respectively.
In Table 7, LLMs exhibit a strong capability in identifying frequent subgraphs, with GPT-4 and O1-mini showing impressive performance. However, LLMs are prone to detect simpler patterns rather than more complex ones in the dataset. In Figure 5, we illustrate the accuracy when different models extract the defined patterns in the datasets. Although various patterns appear with similar probability across different datasets, LLMs predominantly identify triangles. The house pattern, a combination of a triangle and square, is rarely recognized.
@@ 5.3 Discriminative Pattern Learning
The results in Table 8 demonstrate that most LLMs successfully perform graph classification tasks, particularly when using the edge list format. Interestingly, a high D.P. score does not always linearly correspond to high accuracy. For instance, despite having a D.P. score of 0.08, O1-mini still achieves a classification accuracy of 0.72.
@@ 6 Evaluation on Real-world Graphs
In this section, we gather a range of real-world datasets across diverse domains of graphs, including molecules, social networks, bioinformatics, and computer vision. Different from synthetic dataset, the real-world graphs have node attributes and potential noise in the graph structures, which makes the graph pattern tasks more challenging. More details of datasets and prompts are provided in the appendix B. In this evaluation, we focus on pattern detection and discriminative pattern learning.
@@ 7 Discussions
In this section, we first provide our theoretical intuitions from the perspective of expressiveness and then summarize key insights from our observations.
@@ 7.1 Theoretical Intuitions
Here we discuss the theoretical intuitions of LLMs’ capabilities in graph pattern recognition from an expressiveness standpoint, often used in analyzing Transformers (Feng et al., 2024; Li et al., 2024; de Luca & Fountoulakis, 2024). We demonstrate that Transformers, given graph inputs, can be configured to perform graph pattern tasks. The LOCAL framework Angluin (1980) is a distributed computing paradigm, implemented by the message passing among the nodes. Many graph pattern algorithms such as pattern detection can be efficiently implemented by LOCAL  (Drucker et al., 2014; Korhonen & Rybicki, 2017). Intuitively if Transformers can simulate any LOCAL algorithms, it has the capability for graph pattern tasks. We build the following theorem for this intuition.
@@
The detailed proof is given in Appendix A. The proof intuition is that the attention mechanism is a message-passing between the tokens and each token is executed in a parallel manner. Consequently, this mechanism enables Transformers to simulate the message passing between nodes. The above theorem suggests the existence of such weights.
@@ 7.2 Empirical Insights
LLMs use diverse algorithms for one task, and the performance varies due to their execution ability. We provide two observations: (1) We manually reviewed most of the outputs generated by LLMs in graph mining tasks, and summarized the algorithms used by LLMs in Appendix H. Our analysis reveals that different LLMs utilize diverse algorithms to solve the same problem. For instance, more than eight algorithms are used for pattern detection tasks (Section 3.3). (2) Due to the internal flaws of LLMs, these algorithms, although logically correct, will have different performance. In the graph isomorphic mapping task (Section 4.1), a common algorithm starts by counting node degrees and then mapping nodes. O1-mini uses this approach for 89% of the data but achieves only 30% accuracy due to errors in degree counting. In contrast, Claude applies degree counting to only 23% of the data, relying primarily on a direct edge-matching algorithm for the rest. This alternative strategy enables Claude to achieve an impressive 96% accuracy.
Input format that aligns with the pretrained knowledge improves the performance. Our experiments demonstrate that LLMs have basic knowledge about graph patterns and they help LLMs in graph pattern tasks. First, LLMs are pre-trained on extensive internet datasets where graph patterns are often described using specific terminologies. This exposure helps LLMs understand these terms. Comparing the results in Section 3.3 with those in Section 4.3, we observe that terminology-based graph pattern detection generally outperforms topology-based detection. This suggests that LLMs leverage their internal knowledge to enhance performance when provided with terminology as input. Second, the pretrained knowledge will influence the strategies employed by LLMs, and the graph input format that aligns with the strategies will improve the performance. For example, in the case of discriminative pattern learning (Section 5.3), the algorithms used by LLMs often rely on comparing corresponding edges in two graphs. In this scenario, the edge list format typically leads to better performance than the adjacency list format. Conversely, in k-core detection (Section 5.1), the algorithms require counting node degrees, so the edge list is inferior to the adjacency list.
@@ 8 Related work
Large Language Models for Graphs. Here, we introduce previous work using LLMs based on the types of graph tasks they focused on. The first category is to solve graph algorithm tasks, such as connectivity, degree counting, cycle check, shortest path, and Hamilton path. NLGraph (Wang et al., 2024) and Talk like a graph (Fatemi et al., 2023) translated graph structures into natural language, and leveraged techniques like few-shot prompting or chain-of-thought (CoT) to feed these descriptions directly into LLMs for inference. Since exact computational algorithms can be employed to obtain an accurate reasoning process and results, GraphInstruct (Luo et al., 2024) built the instruction-tuning dataset and fine-tuned LLMs by Lora. Moreover, Graphwiz (Chen et al., 2024a) extended the scale and diversity of the dataset and applied Direct Preference Optimization (DPO) loss to enhance the performance. Recently, GraphToken (Perozzi et al., 2024) utilized a GNN-like encoder to process the input graph and train this encoder with a frozen LLM.
@@
Graph Pattern Discovery. Graph patterns (also known as substructures, graphlets, motifs and subgraphs) have been extensively studied before the era of LLMs due to their crucial role in real-world applications, particularly in chemistry (Murray & Rees, 2009), biology (Hu et al., 2005), and social science (Wu et al., 2022). Numerous algorithms have been proposed for pattern discovery. For example, the algorithms designed to identify and count dense pre-defined patterns, such as k-core, k-cycle, and k-clique, were introduced in  Milo et al. (2002); Zhang & Parthasarathy (2012). Frequent subgraph mining algorithms were developed to efficiently search for possible patterns (Yan et al., 2008; Elseidy et al., 2014), aiming to reduce the search space as much as possible.
In recent years, deep learning, particularly graph neural networks (GNNs), has become prominent in this area. Many studies (Chen et al., 2020; Zhang et al., 2023; Luo et al., 2022) investigated the theoretical guarantees of specific GNNs capable of identifying certain substructures, and developed subgraph-aware GNNs with strong expressiveness to improve the accuracy of graph learning tasks. More recently, some studies (Zheng et al., 2023) have begun exploring the use of LLMs to identify SMILES patterns in molecular data, opening new avenues for research.
@@ 9 Conclusion
Acknowledgments
The work is supported by the Accelerating Foundation Model Research Program (AFMR), Microsoft Research. Xinnan Dai and Jiliang Tang are supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, the Army Research Office (ARO) under grant number W911NF-21-1-0198, Amazon Faculty Award, Meta, Microsoft and SNAP.
Appendix A Proof for Theorem 1
@@
Appendix B Dataset
B.1 Primitive graph patterns
We select primitive graph patterns with varying node counts and edge numbers. First, 3-node patterns are the simplest structures in graphs, so we begin with patterns like the triangle, V-structure (V-S), feedforward loop (FFL), and feedback loop (FBL). In directed graphs, the V-S has two edges, while both the FFL and FBL have three. However, the FFL and FBL differ in the direction of one edge. For 4-node patterns, we select the tailed-triangle (T-triangle), square, and diamond for undirected graphs, and the directed-diamond for directed graphs. Both the T-triangle and square have four edges but differ in their connectivity, while the diamond includes five edges. Finally, we introduce the house pattern, a 5-node structure combining a triangle and a square. The summarization of patterns shown in Table 11.
@@ B.1 Primitive graph patterns
B.2 Synthetic dataset details
@@
B.3 Real-world dataset details